mario's age is 14 years more than 6 times juan's age. The sum of their ages is 35. How old is each person.

Answer by stanbon(75887)

mario's age is 14 years more than 6 times juan's age. The sum of their ages is 35. How old is each person.
------------------
M = 6J + 14
M + J = 35
-----------------------
Substitute for "M" and solve for "J"::
M = 35-J
-----Ch
35-J = 6J + 14
-------------------
7J = 21
Juan = 3 yrs
------
Mario = 35-J = 32 yrs
----------------
Cheers,
Stan H.
-------------